OpenStudy (anonymous):

you have 5/6 cup of flour in your cupboard . a recipe calls for 1/4 cup of flour . how much flour do you have left if you made the bread ?

OpenStudy (anonymous):

5/6-1/4=10/12-3/12=7/12 cup left

OpenStudy (anonymous):